2- Draw the projection of b onto a and compute it from p= ła. a) a=[1], 6-[4] b) a=[tan ® ], v=[sin )
Answer #1

p= ña ; p projection of to onto a wheree ñ = arb ata - 1 -1 = 0 ata = - 1 + 1 = 2 sap= - 0011-01 (1,1)tan o I Sino Woohoo E tan o p= a ña (Projection of t anta a a an o tan² o tan atb= [tano of sino = sino tan ara= [rano of si